OpenStudy (anonymous):

eleventh root of x to the fifth power over eighth root of x to the third power . Write the simplified answer in radical form.

OpenStudy (saifoo.khan):

\[\Huge \frac{\sqrt[11]{x^5}}{\sqrt[8]{x^3}}\]

OpenStudy (saifoo.khan):

Like that>

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thats how it looks like

OpenStudy (saifoo.khan):

ok, so you can write it as:

OpenStudy (saifoo.khan):

\[\Huge \frac{x^\frac{5}{11}}{x^\frac38}\]

OpenStudy (anonymous):

would the answer be \[\sqrt[88]{x} ^{7}\]

OpenStudy (saifoo.khan):

\[\Huge x^{\frac{5}{11}-\frac38}\]

OpenStudy (saifoo.khan):

Yes, you are right.

OpenStudy (anonymous):

thank you very very much :)

OpenStudy (saifoo.khan):

you welcome! :)

OpenStudy (anonymous):

Alright have a nice day and once again thanks for answering my question:)

OpenStudy (saifoo.khan):

Same to you!! No problem!!

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

[: